Smith Enterprises (DAX: SME) Last Trade: Trade Time: Change: Prev Close: Open Bid Ask: ly Target Est: 55.40 $51.63 4:00 PM ET ??? (12.09%) 46.06 47.25 N/A N/A Days Range: 47.22 - 51.96 52wk Range 25.48-60.71 Volume: Avg Vol (3m 629,072 Market Cap: P/E (ttm): EPS (ttm): Div & Yield: 1,419,317 1.63B 23.58 2? 0.20 (0.40%) The last trade time is shown to be at 4:00 PM Eastern Time. The abbreviation ttm stands for trailing twelve months, and the number shown reflects data from the previous 12 months. The abbreviation 3m reflects data from the previous 3 months What exchange does Smith Enterprises trade on? O NASDAQ SME O FTSE DAX O AMEX What is Smiths earnings per share (EPS) for the trailing 12 months? $2.03 O $2.14 $1.83 O $1.91 O $2.19 How much did the price increase today over yesterdays closing price? O $5.57 $5.29 O $5.50 $5.43 O $5.36

5. how heavy was the trading of Smith's stock today relative to the normal level of trading?

relatively heavy trading

average trading

relatively light trading
Answer #1

The stock exchange is :DAX

It is mentioned: (DAX:SME)

EPS shall be calculated from PE ratio

PE=Price/EPS
Price=51.63

PE=23.58

EPS=51.63/23.58=$2.19

Relatively heavy trading as it is more than double of average 3 months trading

Increase=51.63-46.06=$5.57